what is the concentration of a solution with a volume of 3.3ml that contains 12 grams of ammonium sulfite((NN4)2SO3)


1
Expert's answer
2022-04-21T02:01:04-0400

12 / 116.15 = 0.1033 mol of ammonium sulfite

V = 3.3 mL = 0.0033 L of the solution

Molarity = 0.1033 / 0.0033 = 31.3 M